pam is 14 and dad is 37. In how many years will pam's dad be twice as old as she will be?
Answer by ankor@dixie-net.com(22740)

pam is 14 and dad is 37.
In how many years will pam's dad be twice as old as she will be?
:
let y = no. of years for this to be true
:
37 + y = 2(14 + y)
37 + y = 28 + 2y
37 - 28 = 2y - y
9 = y
:
In 9 years Dad will be twice Pam's age
:
:
Confirm this
37 + 9 = 2(14 + 9)
46 = 2(23)